OpenStudy (anonymous):

2 STEP EQUATION 14 - 9x = -8(10 + x)

OpenStudy (anonymous):

First distribute the -8 14-9x=-80-8x add 9x to both sides 14=-80+x add 80 to both sides 94=x

OpenStudy (anonymous):

Thank You So Much!!!